Question:
x(squared)+6x-27=0

Answer by tutorcecilia(2152)

x^2+6x-27=0
(x-3)(x+9)=0 [factor]
x-3=0 [set each factor equal to zero]
x=3
.
x+9=0
x=-9
.
check by plugging either x=3 or x=-9 back into the original equation and solve.